If the Ksp of Ag2CO3 at 25 oC is 8.46e-12 , what is the molar solubility of Ag2CO3 at this temperature?

a) At this temperature the molar solubility of Ag2CO3 is 6.05e-34 mol/L.
b) At this temperature the molar solubility of Ag2CO3 is 2.04e-04 mol/L.
c) At this temperature the molar solubility of Ag2CO3 is 2.91e-06 mol/L.
d) At this temperature the molar solubility of Ag2CO3 is 1.28e-04 mol/L.
e) At this temperature the molar solubility of Ag2CO3 is 2.42e-33 mol/L.
Answer #1

At equilibrium:

Ag2CO3 <----> 2 Ag+ + CO32-

   2s s

Ksp = [Ag+]^2[CO32-]

8.46*10^-12=(2s)^2*(s)

8.46*10^-12= 4(s)^3

s = 1.284*10^-4 M

Answer: d